A has 3 tickets in a lottery containing 3 prizes and 9 blanks;B has 2 tickets in lottery containing 2 prizes and 6 blanks. copare their chances of succes?

Grade:11

1 Answers

Aditi Yandagoudar
7 years ago
For A, there can be three cases :-1st case: WWW = 3C3 × 9C0 =92nd case:WWL= 3C2 × 9C1 =273rd case : WLL= 3C1 * 9C2=108P (A) = 9+27+108/ 12C3 =34/55Similarly P (B) = ( 2C2 × 6C0 + 2C1 ×6C1 ) / 8C2 = 13/28 P(A):P(B) = 34/55:13/28 = 952:715